# Pulling Back the Curtain on ReLU Networks

### v3

@@
Since any ReLU network is piecewise affine, its hidden units can be characterized by their pullbacks through the active subnetwork, i.e., by their gradients (up to bias terms). However, gradients of deeper neurons are notoriously misaligned, which obscures the network’s internal representations. We posit that models do align gradients with data, yet this is concealed by the intrinsic noise of the ReLU hard gating. We validate this intuition by applying soft gating in the backward pass only, reducing the local impact of weakly excited neurons. The resulting modified gradients, which we call excitation pullbacks, exhibit striking perceptual alignment on a number of ImageNet-pretrained architectures, while the rudimentary pixel-space gradient ascent quickly produces easily interpretable input- and target-specific features. Inspired by these findings, we formulate the path stability hypothesis, claiming that the binary activation patterns largely stabilize during training and get encoded in the pre-activation distribution of the final model. When true, excitation pullbacks become aligned with the gradients of a kernel machine that mainly determines the network’s decision. This provides a theoretical justification for the apparent faithfulness of the feature attributions based on these pullbacks, potentially even leading to mechanistic interpretability of deeper models. Incidentally, we give a possible explanation for the effectiveness of Batch Normalization and Deep Features, together with a novel perspective on the network’s internal memory and generalization properties. We release the code and an interactive app for easier exploration of the excitation pullbacks.
@@ 1 Introduction
Understanding how neural networks organize their internal representations remains a longstanding and fundamental challenge in deep learning. In this paper we report the results of a less common line of research, where our primary goal was not incremental leaderboard improvement, but rather breaking through conceptual barriers. We’ve published111github.com/314-Foundation/ResearchLog the weekly research log to facilitate the discussion on this style of scientific inquiry. While we reduce the problem space to ReLU convolutional networks for vision tasks, our methods likely generalize to other activations with similar gating behavior, and potentially offer insights into a broader range of architectures and modalities.
@@
The striking perceptual alignment of excitation pullbacks strongly suggests a deeper cause. Our primary concern was to validate theoretically that the produced explanations are reasonably faithful to the network’s decision process. Building on our previous investigations we developed a formal framework that allows us to analyze the pullbacks from the path perspective, where path is any sequence of neurons selected from the consecutive layers. In particular, we observe that ReLU networks correspond to linear models in the path space (the Hilbert space of formal linear combinations of all paths), under a feature map given by the tensor product of the binary activation vectors (encoding the on/off states of neurons) of subsequent layers (see Section 3 for details). This made us suspect that these feature maps might largely stabilize during training in the form of a higher-rank tensor field, which we call a path tensor feature (see Hypothesis 1). If that’s indeed the case, then the network behaves like a piecewise linear kernel machine with excitation pullbacks approximating directionally it’s gradient, so that the feature attributions they produce can be considered faithful to the network’s internal circuitry. This would translate to excitation pullbacks constituting a mechanistic interpretation of deep models in a style reminiscent of the “neural circuits” program (Cammarata et al., 2020) - but considerably streamlined by the pullback computation.
Incidentally, the above considerations provide a possible explanation for the widespread effectiveness of Batch Normalization (Ioffe and Szegedy, 2015) in accelerating training, as normalizing neural activations significantly speeds up the stabilization of activation patterns and, consequently, the emergence of the path tensor feature and the global linear structure of the network’s internal representations, see Section 5.3. Furthermore, the success of Deep Features (Zhang et al., 2018) as a perceptual metric could be attributed to its kernel structure and the close relation to the path tensor feature (see Section 5.2). Interestingly, path tensor features allow us to view neural networks as associative memory systems, with excitation pullbacks as their contents (Section 5.4).
Our work demonstrates that standard pretrained networks do in fact align gradients with data, and highlights how one can extract interpretable representations of hidden neurons. Furthermore, the proposed formal framework sheds light on the internal organization of neural networks, and even offers a potential explanation for their generalization capabilities: if the stability condition expressed in Hypothesis 1 is satisfied, then the network’s decisions can be well approximated by a kernel machine whose gradient is feasibly accessible by the excitation pullback. The striking perceptual alignment observed in our experiments strongly suggests that this may indeed be the case, at least to some extent.
@@
Section 2 introduces necessary notions and notational conventions; in particular, the gating function and gating-induced pullbacks; Section 3 defines tensor field pullbacks and shows that ReLU networks correspond to linear maps in the path space; Section 4 introduces the excitation pullbacks and discusses the inherent gradient noise; Section 5 discusses the path stability hypothesis and its implications; Section 6 details the pullback visualization method used in the paper; Section 7 describes the related works and discusses the limitations of presented approach.
We release the code222github.com/314-Foundation/ExcitationPullbacks and the interactive demo app333 huggingface.co/spaces/msat/ExcitationPullbacks to make the exploration of excitation pullbacks easily accessible.
@@ 3.2 Tensor field pullbacks
$\tilde{\Lambda}(x)\cong\bigotimes_{\ell=1}^{L}\lambda_{\ell}(x)\in\bigotimes_{\ell=1}^{L}\mathbb{R}^{d_{\ell}}\cong\mathbb{R}^{P_{1}}$ (13)
@@ 3.3 Path space and the linear nature of ReLU networks
$\Omega((y,\theta)):=(\omega_{p})_{p\in\mathcal{P}_{0}},\quad\text{where }\omega_{p}:=y[p_{L}]\prod_{l=1}^{L}W_{\ell}[p_{\ell},p_{{\ell-1}}],\quad p=(p_{0},p_{1},\dots,p_{L})$ (16)
@@
$\phi_{\tau}(x):=(\tau_{p}\cdot x[p_{0}])_{p\in\mathcal{P}_{0}}\cong\tau\otimes x$ (17)
@@
$f_{\tau}(x)=\sum_{p\in\mathcal{P}_{0}}(\phi_{\tau}(x))_{p}\cdot\omega_{p}=\langle\phi_{\tau}(x),\omega\rangle$ (18)
Observe that the dimension of $\mathbb{R}^{P_{0}}$ scales multiplicatively with the network’s depth and therefore, in deeper networks, it is orders of magnitude larger than the number of network parameters (which scales additively with depth). In a sense this means that neural networks, generally considered to be overparameterised models, are actually significantly underparameterised from the path space point of view.
@@
Note 3.3 The equation 19 explicitly disentangles network’s weights from the feature function, as, by definition, $\omega$ does not depend on $\tau$ and $\phi_{\tau}$ does not depend on the network parameters $(y,\theta)$. This disentangled view suggests that most of the network’s expressive power is due to the highly non-linear gating function $G$, as the network acts linearly (via $\omega\in\mathbb{R}^{P_{0}}$) on inputs transformed by $\phi_{\tilde{G}}$. Moreover, by Equation 19, the action of $\omega$ on $\phi_{\tilde{G}}(x)$ can be directly visualised in the input space via the action of the pullback $v_{G}$ on $x$. Even though $\phi_{\tilde{G}^{(t)}}$ does change during training, it is significantly more stable than the direct mapping $M_{G}^{(t)}:x\mapsto x_{L}^{(t)}\in\mathbb{R}^{d_{L}}$, hinting at the importance of this representation.
@@ 4.2 Introducing excitation pullbacks
The considerations from previous section suggest that we should look beyond the hard activation gates $G(x)$ to properly pull the action of $y^{\top}$ back to $\mathbb{R}^{d_{0}}$ - in a way that is robust to hyper-local noise. A natural idea is to perform soft gating, i.e. compose the pre-activations $z_{\ell}$ with some sigmoid-like step function $\sigma_{\ell}:\mathbb{R}^{d_{\ell}}\to[0,1]^{d_{\ell}}$:
$\Gamma:\mathbb{R}^{d_{0}}\to\prod_{\ell=1}^{L}[0,1]^{d_{\ell}},\quad\Gamma:=(\gamma_{1},\dots,\gamma_{L}),\quad\gamma_{\ell}:=\sigma_{\ell}\circ z_{\ell}:\mathbb{R}^{d_{0}}\to[0,1]^{d_{\ell}}$ (21)
In general, $\sigma_{\ell}$ is neuron-specific, since the pre-activation distribution may be different for each neuron. In practice, thanks to the BatchNorm layers, a shared $\sigma$ across all neurons yields good empirical results (even though BatchNorm still differentiates between the neuron distributions via it’s affine parameters). In the paper we set the global element-wise $\sigma_{\ell}=\sigma=\operatorname{sigmoid(\frac{z}{0.3})}:\mathbb{R}\to[0,1]$, but other options (e.g. rescaled softsign or Normal CFD) lead to very similar results - as long as the general sigmoid shape of the global $\sigma$ is preserved, see Section 6.2.
Because $\sigma_{\ell}$ can be understood as measuring the excitation of a neuron we may accordingly call $\Gamma$ the excitation function, it’s corresponding pullback $v_{\Gamma}$ the excitation pullback, and it’s induced tensor field $\tilde{\Gamma}$ the excitation tensor field. Note that in order to compute $\Gamma$ we still need to use the original gates $G$ in the forward pass to compute the right pre-activations.
Note 4.2 The tensor field $\tilde{\Gamma}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{P_{1}}$ (see Equation 12) suppresses barely active paths, i.e. those passing through weakly activated neurons; and partially recovers paths that are nearly active (i.e. those interrupted by a single weak neuron, but otherwise strong). This makes $v_{\Gamma}$ less sensitive to hyper-local noise across linear regions when compared with $v_{G}$ and probably more aligned with the coarser structure of the decision boundary. In a sense, $v_{\Gamma}$ is dominated by atoms corresponding to highly excited paths, as it quickly suppresses the less-excited paths (see Equation 10).
@@ 5.1 The path stability hypothesis
Because of the relative stability of $\phi_{\tilde{G}^{(t)}}$ (Note 3.3), it’s tempting to assume that the gating function $G^{(t)}$ gets fixed early on during training. We consider this assumption too optimistic as the gates probably do change simply because of the gradient noise we’ve hinted at before. Notice, however, that $\tilde{G}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{P_{1}}$ maps inputs to binary rank-1 tensors, meaning that it has limited path-filtering capability (Note 3.1). But perhaps there are more refined atom filters (corresponding to higher-rank binary tensors) that do get fixed early on during training? We motivate this assumption below.
@@
Earlier work has demonstrated that only a small subnetwork within a larger model is responsible for most of its performance. In particular, weight pruning techniques demonstrate that many weights in trained networks can be removed without significantly degrading accuracy, implying the presence of sparse, efficient subnetworks within dense architectures (Han et al., 2015). Moreover, the Lottery Ticket Hypothesis (Frankle and Carbin, 2019) posits that such performant subnetworks often emerge early during training and can be “rewound” to their initial weights to train successfully in isolation. Furthermore, in (Lakshminarayanan and Singh, 2021) the authors define Neural Path Features (NPF), which are essentially our gating functions (Equation 3), and argue that “almost all the information learnt by a DNN with ReLU activations is stored in the gates”. Specifically, they introduce Deep Gated Networks (DGN) and provide empirical evidence that gate adaptation is the key to generalization, most notably, that “the winning lottery is in the gating pattern”  (Lakshminarayanan and Singh, 2020). Together, these observations motivate the idea that the gating patterns in ReLU networks may approximately stabilize early in training, and thus may carry semantically meaningful structure long before convergence.
Recall that $f^{(t)}(x)=\langle\omega^{(t)},\phi_{\tilde{G}^{(t)}}(x)\rangle$ and $f^{(t)}_{\tau}(x)=\langle\omega^{(t)},\phi_{\tau}(x)\rangle$. Let $X=(x_{i})_{i}\subset\mathbb{R}^{d_{0}}$ be the training dataset and $f(X)=(f(x_{i}))_{i}$ be the vector of all the values of $f$. We say that two vectors $a,b\in\mathbb{R}^{n}$ are strongly positively correlated if the Pearson correlation coefficient $\rho(a,b)$ is high (e.g., greater than 0.9, although the exact threshold value is not important). We denote this relation as $a\approx_{+}b$. We now state our core hypothesis:
Hypothesis 1 (Early emergence of stable paths).
@@
There exists a piecewise constant binary tensor field $\pi:\mathbb{R}^{d_{0}}\to\{0,1\}^{{P_{1}}}$ and an early training time $t_{\pi}\ll T$ such that for every $t>t_{\pi}$ we have $f^{(t)}(X)\approx_{+}f^{(t)}_{\pi}(X)$.
Additionally, we postulate that the network encodes the tensor field $\pi$ in its pre-activation function $Z$. That is, there exists a choice of $(\sigma_{\ell})_{\ell}$ (as in Equation 21) such that for every $x\in X$ we have $\tilde{\Gamma}(x)\approx_{+}\pi(x)$.
@@
We refer to the above assumption as the path stability condition and to any such $\pi$ as path tensor feature.
Intuition:
@@
The hypothesis states that, even though the gating tensors $\tilde{G}^{(t)}(x)$ change during training, there exists a time-independent gating tensor $\pi(x)$ that, starting from time $t_{\pi}$, largely determines network’s decision $f^{(t)}(x)$ at $x$. In this view, the variability of tensors $\tilde{G}^{(t)}(x)$ can be attributed to their limited expressive power as rank-1 tensors; the network, trying to approximate $\pi(x)$, learns it implicitly in it’s pre-activation pattern $Z(x)$ in the form of the set of highly-excited paths. The latter can be feasibly directionally approximated by the rank-1 non-binary tensor $\tilde{\Gamma}(x)$.

### v4

@@
Since any ReLU network is piecewise affine, its hidden units can be characterized by their pullbacks through the active subnetwork, i.e., by their gradients (up to bias terms). However, gradients of deeper neurons are notoriously misaligned, which obscures the network’s internal representations. We posit that models do align gradients with data, yet this is concealed by the intrinsic noise of the ReLU hard gating. We validate this intuition by applying soft gating in the backward pass only, reducing the local impact of weakly excited neurons. The resulting modified gradients, which we call excitation pullbacks, exhibit striking perceptual alignment on a number of ImageNet-pretrained architectures, while the rudimentary pixel-space gradient ascent quickly produces easily interpretable input- and target-specific features. Inspired by these findings, we formulate the path stability hypothesis, claiming that the binary activation patterns largely stabilize during training and get encoded in the pre-activation distribution of the final model. When true, excitation pullbacks become aligned with the gradients of a kernel machine that mainly determines the network’s decision. This provides a theoretical justification for the apparent faithfulness of the feature attributions based on excitation pullbacks, potentially even leading to mechanistic interpretability of deep models. Incidentally, we give a possible explanation for the effectiveness of Batch Normalization and Deep Features, together with a novel perspective on the network’s internal memory and generalization properties. We release the code and an interactive app for easier exploration of the excitation pullbacks.
@@ 1 Introduction
Understanding how neural networks organize their internal representations remains a longstanding and fundamental challenge in deep learning. In this paper we report the results of a less common line of research, where our primary goal was not incremental leaderboard improvement, but rather breaking through conceptual barriers. We’ve published [^1] https://github.com/314-Foundation/ResearchLog the weekly research log to facilitate the discussion on this style of scientific inquiry. While we reduce the problem space to ReLU convolutional networks for vision tasks, our methods likely generalize to other activations with similar gating behavior, and potentially offer insights into a broader range of architectures and modalities.
@@
The striking perceptual alignment of excitation pullbacks strongly suggests a deeper cause. Our primary concern was to motivate theoretically that the produced explanations are reasonably faithful to the network’s decision process, as this is the main unsolved issue in the attribution research (Deng et al., 2025). Building on our investigations we developed a formal framework that allows us to analyze the pullbacks from the path perspective, where path is any sequence of neurons selected from the consecutive layers. In particular, we’ve observed that ReLU networks correspond to linear models in the path space (the Hilbert space of formal linear combinations of all paths), under a feature map given by the tensor product of the binary activation vectors (encoding the on/off states of neurons) of subsequent layers (see Section 3 for details). This made us suspect that these feature maps might largely stabilize during training in the form of a higher-rank tensor field, which we call a path tensor feature (see Hypothesis 1). If that’s indeed the case, then the network behaves like a piecewise linear kernel machine with excitation pullbacks approximating directionally it’s gradient, so that the feature attributions they produce can be considered faithful to the network’s internal circuitry. This would translate to excitation pullbacks constituting a mechanistic interpretation of deep models in a style reminiscent of the “neural circuits” program (Cammarata et al., 2020), but considerably streamlined by the pullback computation.
We view our hypothesis primarily as a lens through which neural networks can be studied, rather than a statement that is universally valid. We expect it to hold only partially, but nonetheless consider its formulation useful for guiding further research. In particular, it inspires a possible explanation for the widespread effectiveness of Batch Normalization (Section 5.3) and the success of Deep Features as a perceptual metric (Section 5.2). Even more interestingly, it allows us to view classification networks as associative memory systems, with excitation pullbacks as their contents (Section 5.4).
Our work demonstrates that standard pretrained networks do seem to align gradients with data, and highlights how one can extract interpretable and arguably faithful representations of hidden neurons. Furthermore, the proposed formal framework sheds light on the internal organization of neural networks, and even offers a potential explanation for their generalization capabilities: if the path stability condition expressed in Hypothesis 1 is satisfied, then the network’s decisions can be well approximated by a kernel machine whose gradient is feasibly computable by the excitation pullback. The striking perceptual alignment observed in our experiments strongly suggests that this may indeed be the case, at least to some extent. While our formal contributions remain preliminary and partly speculative, we see them as pointing toward a potentially rich avenue for future work.
@@
Section 2 introduces necessary notions and notational conventions; in particular, the gating function and gating-induced pullbacks; Section 3 defines tensor field pullbacks and shows that ReLU networks correspond to linear maps in the path space; Section 4 introduces the excitation pullbacks and discusses the inherent gradient noise; Section 5 presents the path stability hypothesis and its implications; Section 6 details the pullback visualization method used in the paper; Section 7 describes the related works and discusses the limitations of presented approach.
We release the code [^2] https://github.com/314-Foundation/ExcitationPullbacks and the interactive demo app [^3] https://huggingface.co/spaces/msat/ExcitationPullbacks to make the exploration of excitation pullbacks easily accessible.
@@ 3.2 Tensor field pullbacks
$\tilde{\Lambda}(x)=\bigotimes_{\ell=1}^{L}\lambda_{\ell}(x)\in\bigotimes_{\ell=1}^{L}\mathbb{R}^{d_{\ell}}\cong\mathbb{R}^{P_{1}}$ (13)
@@ 3.3 Path space and the linear nature of ReLU networks
$\Omega((y,\theta)):=\omega=(\omega_{p})_{p\in\mathcal{P}_{0}},\quad\text{where }\omega_{p}:=y[p_{L}]\prod_{l=1}^{L}W_{\ell}[p_{\ell},p_{{\ell-1}}],\quad p=(p_{0},p_{1},\dots,p_{L})$ (16)
@@
$\phi_{\tau}(x):=(\tau_{p}(x)\cdot x[p_{0}])_{p\in\mathcal{P}_{0}}=\tau(x)\otimes x$ (17)
@@
$f_{\tau}(x)=\sum_{p\in\mathcal{P}_{0}}(\phi_{\tau}(x))_{p}\cdot\omega_{p}=\langle\phi_{\tau}(x),\omega\rangle=\langle\tau(x)\otimes x,\omega\rangle$ (18)
Observe that the dimension of $\mathbb{R}^{P_{0}}$ scales multiplicatively with the network’s depth and therefore, in deeper networks, it is orders of magnitude larger than the number of network parameters (which scales additively with depth). In a sense, this let’s us consider neural networks as underparameterised models on the path space.
@@
Note 3.3 Equation 18 explicitly disentangles network’s weights from the feature function, as, by definition, $\omega$ does not depend on $\tau$, and $\phi_{\tau}$ does not depend on the network parameters $(y,\theta)$. This disentangled view suggests that most of the network’s expressive power is due to the highly non-linear gating function $G$, as the network acts linearly (via $\omega\in\mathbb{R}^{P_{0}}$) on inputs transformed by $\phi_{\tilde{G}}$. Moreover, by Equation 19, the action of $\omega$ on $\phi_{\tilde{G}}(x)$ can be directly visualised in the input space via the action of the pullback $v_{G}$ on $x$. Even though $\phi_{\tilde{G}^{(t)}}$ does change during training, it is significantly more stable than the direct mapping $M_{G}^{(t)}:x\mapsto x_{L}^{(t)}\in\mathbb{R}^{d_{L}}$, hinting at the importance of this representation.
@@ 4.2 Introducing excitation pullbacks
The considerations from previous section suggest that we should look beyond the hard activation gates $G(x)$ to properly pull the action of $y^{\top}$ back to $\mathbb{R}^{d_{0}}$ - in a way that is robust to hyper-local noise. A natural idea is to perform soft gating, i.e. compose the pre-activations $z_{\ell}$ with some sigmoid-like step functions $\sigma_{\ell,i}:\mathbb{R}\to[0,1]$ applied element-wise:
$\Gamma:\mathbb{R}^{d_{0}}\to\prod_{\ell=1}^{L}[0,1]^{d_{\ell}},\quad\Gamma:=(\gamma_{1},\dots,\gamma_{L}),\quad\gamma_{\ell}:=\sigma_{\ell}\circ z_{\ell}:\mathbb{R}^{d_{0}}\to[0,1]^{d_{\ell}},\quad\sigma_{\ell}(z_{\ell}):=\bigl(\sigma_{\ell,i}(z_{\ell}[i])\bigr)_{i=1}^{d_{\ell}}$ (21)
In general, we use neuron-specific functions $\sigma_{\ell,i}$, since the pre-activation distribution may be different for each neuron. In practice, thanks to the BatchNorm layers, a shared $\sigma$ across all neurons yields good empirical results (even though BatchNorm still differentiates between the neuron distributions via it’s affine parameters). In the paper we use a global element-wise $\sigma_{\ell,i}=\sigma=\operatorname{sigmoid(\frac{z}{0.3})}$, but other options (e.g. rescaled softsign or Normal CFD) lead to very similar results - as long as the general sigmoid shape of the global $\sigma$ is preserved, see Section 6.2.
Because $\sigma_{\ell,i}$ can be understood as measuring the excitation of a neuron we accordingly call $\Gamma$ the excitation function, it’s corresponding pullback $v_{\Gamma}$ the excitation pullback, and it’s induced tensor field $\tilde{\Gamma}$ the excitation tensor field. Note that in order to compute $\Gamma$ we still need to use the original gates $G$ in the forward pass to find the right pre-activations.
Note 4.2 The tensor field $\tilde{\Gamma}:\mathbb{R}^{d_{0}}\to\mathbb{R}^{P_{1}}$ (see Equation 12) suppresses barely active paths, i.e. those passing through weakly activated neurons; and partially recovers paths that are nearly active (i.e. those interrupted by a single weak neuron, but otherwise strong). This makes $v_{\Gamma}$ less sensitive to hyper-local noise across linear regions when compared with $v_{G}$ and more likely to be better aligned with the coarser structure of the decision boundary. In a sense, $v_{\Gamma}$ is dominated by atoms corresponding to highly excited paths, as it quickly suppresses the less-excited paths (see Equation 10).
@@ 5 On the path stability and its potential significance
This section should be regarded as exploratory. Our aim here is not to establish formal results, but rather to outline conjectures and highlight promising research avenues. To this end, we don’t shy away from semi-formal reasoning and intuitive arguments, leaving rigorous analysis for future work.
@@ 5.1 The path stability hypothesis
Earlier work has demonstrated that only a small subnetwork within a larger model is responsible for most of its performance. In particular, weight pruning techniques demonstrate that many weights in trained networks can be removed without significantly degrading accuracy, implying the presence of sparse, efficient subnetworks within dense architectures (Han et al., 2015). Moreover, the Lottery Ticket Hypothesis (Frankle and Carbin, 2019) posits that such performant subnetworks often emerge early during training and can be “rewound” to their initial weights to train successfully in isolation. Furthermore, in (Lakshminarayanan and Singh, 2021) the authors define Neural Path Features (NPF), which are essentially our gating functions (Equation 3), and argue that “almost all the information learnt by a DNN with ReLU activations is stored in the gates”. Specifically, they introduce Deep Gated Networks (DGN) and provide empirical evidence that gate adaptation is the key to generalization, most notably, that “the winning lottery is in the gating pattern”  (Lakshminarayanan and Singh, 2020). Together, these observations motivate the idea that the gating patterns in ReLU networks may approximately stabilize early in training, and thus may carry semantically meaningful structure long before convergence. That being said we consider it too naive to assume that the gating tensors $\tilde{G}^{(t)}$ get fixed, because of their limited path-filtering capability (Note 3.1) as rank-1 binary tensors and the gradient noise we’ve hinted at before. However, paths may stabilize as higher-rank tensors.
Let $\omega^{(t)}=\Omega((y^{(t)},\theta^{(t)}))\in\mathbb{R}^{P_{0}}$ encode the network’s parameters at time $t$. We’ve already established the following:
$f^{(t)}_{\tau}(x)$ $=\langle\omega^{(t)},\phi_{\tau}(x)\rangle=\langle\omega^{(t)},\tau(x)\otimes x\rangle$ (22)
$f^{(t)}(x)$ $=\langle\omega^{(t)},\phi_{\tilde{G}^{(t)}}(x)\rangle=\langle\omega^{(t)},\tilde{G}^{(t)}(x)\otimes x\rangle\quad\text{(substituting $\tilde{G}^{(t)}(x)$ for $\tau$)}$ (23)
Let $X=(x_{i})_{i}\subset\mathbb{R}^{d_{0}}$ be the training dataset and $f_{\tau}(X)=(f_{\tau}(x_{i}))_{i}$ be the vector of all the values of $f_{\tau}$. We introduce three notions of similarity between vectors. We’ll say that the score is high if it exceeds some fixed value in $[0.5,1]$, understanding that the exact threshold is context-dependent. For $a,b\in\mathbb{R}^{n}$, we say that they are strongly positively correlated, denoted $a\sim_{corr}b$, if the Pearson correlation coefficient $\rho(a,b)$ is high. We say that $a$ and $b$ are approximately proportional, denoted $a\sim_{prop}b$, if there exists $\mu>0$ such that $a\sim\mu b$. Finally, for binary vectors $a,b\in\{0,1\}^{n}$, we say that they are consistent, denoted $a\sim_{F_{1}}b$, if their $F_{1}$-score is high, where $F_{1}(a,b)=\frac{2\sum_{i=1}^{n}a_{i}b_{i}}{\sum_{i=1}^{n}a_{i}+\sum_{i=1}^{n}b_{i}}$ is a standard measure of similarity for binary vectors. We now state our core hypothesis:
Hypothesis 1 (Early stabilization of paths).
@@
There exists a piecewise constant binary tensor field $\pi:\mathbb{R}^{d_{0}}\to\{0,1\}^{{P_{1}}}$ and an early training time $t_{\pi}\ll T$ such that $\forall t>t_{\pi},\;f^{(t)}(X)\sim_{corr}f^{(t)}_{\pi}(X)\;\wedge\;\forall x\in X,\;\tilde{G}^{(t)}(x)\sim_{F_{1}}\pi(x)$. Additionally, tensor field $\pi$ can be chosen so that there exists a function $\Gamma$, as in Equation 21, satisfying $\forall_{x\in X}~\tilde{\Gamma}(x)\sim_{prop}\pi(x)$.
@@
We refer to the above assumption as the path stability condition and to any such $\pi$ as a path tensor feature.
Motivation:
@@
The hypothesis formalizes the intuition that, even though the gating tensors $\tilde{G}^{(t)}(x)$ do change during training, for most $x\in X$ they actually oscillate around a core gating tensor $\pi(x)$ which determines the network’s decision $f^{(t)}(x)$ for $t>t_{\pi}$. In this view, the variability of the tensors $\tilde{G}^{(t)}(x)$ can be attributed to their limited expressive power as rank-1 binary approximations of $\pi(x)$. Additionally, we claim that $\pi(x)$ can be largely recovered as the set of those paths $p$ that are relatively highly excited, as measured by the coordinates $\tilde{\Gamma}_{p}(x)$ of the tensor $\tilde{\Gamma}(x)\in[0,1]^{\mathcal{P}_{1}}$.
Note 5.1 The path tensor feature always exists if $t_{\pi}$ is allowed to be close to $T$, the gist of our assumption being that it forms long before convergence.